1. How many lbs, of cream containing 12 percent butter must be added to 1800lbs,
of milk containing 2 percent butterfat to obtain milk having 3 percent butter?
2. How much of a 40 percent solution of alcohol and how much of an 80 percent
solution should be mixed to give 40 gallons of a 50 percent solution?

Answer by checkley77(12844)

.12x+.02*1,800=.03(1,800+x)
.12x+36=54+.03x
.12x-.03x=54-36
.09x=18
x=18/.09
x=200 pounds of 12% cream is used.
Proof:
.12*200+.02*1,800=.03(1,800+200)
24+36=.03*2,000
60=60